OpenStudy (anonymous):

Suppose a ball moves horizontally at a constant velocity v. Meanwhile, a second ball moves with constant acceleration a along a line oriented at an angle theta from the vertical. At time 0, the first ball is a distance h directly above the second ball, and the second ball is instantaneously at rest.What angle theta would result in a collision at some later time t? The answer should involve the known values v, h, and a (Let us assume that v, h, and a are all positive).This collision implies that the positions (x and y coordinates) of both balls are the same at time t. This gives us two equation

OpenStudy (anonymous):

one for the x coordinate and one for the y coordinate: (1/2)at2sin(theta) = v*t h = (1/2)at2cos(theta)

OpenStudy (anonymous):

Suppose a ball moves horizontally at a constant velocity v. Meanwhile, a second ball moves with constant acceleration a along a line oriented at an angle theta from the vertical. At time 0, the first ball is a distance h directly above the second ball, and the second ball is instantaneously at rest. What angle theta would result in a collision at some later time t? The answer should involve the known values v, h, and a (Let us assume that v, h, and a are all positive). This collision implies that the positions (x and y coordinates) of both balls are the same at time t. This gives us two equations, one for the x coordinate and one for the y coordinate: (1/2)at2sin(theta) = v*t h = (1/2)at2cos(theta)